if the density of methanol is 0.793 kg/l, what is its volume needed for making 2.5l of its 2.5 M solution?

Grade:12th pass

2 Answers

Aarti Gupta
9 years ago

Molar mass of methanol = 32 g mol–1
= 0.032 kg mol–1 (converting to Kg)
Molarity of methanol solution=0.793 Kg L-1/ 0.032 Kg mol -1

= 24.78 mol L–1
( As density = mass/volume)
Applying,
M1V1 = M2V2

(24.78 mol L–1) V1= (2.5 L) (0.25 mol L–1)V1
V1 = 0.0252 L
or,
= 25.22 ml
Thanks & Regards
Aarti Gupta
askiitians Faculty
Manas Satish Bedmutha
9 years ago
No. of moles required = M x V = 2.5 x 2.5 = 6.25 mol Now, moles = wt. / molar mass Therefore, wt. taken = 6.25 x 32 gm = 200 gm = 0.2 kg Using volume = mass / density, Vol= 0.2 / 0.793 = 0.2522 L = 252.2 mL